OpenStudy (anonymous):

Simplify. 5/3x^2 + 9/3x Choices: A. 14/3x B. 5x+9/3x^2 C. 14/9x^3 D. 5+9x/3x^2

OpenStudy (anonymous):

You can multiply the second term by x/x, giving you 9x/3x^2. That way you have like terms.

OpenStudy (anonymous):

I just want the answer. .-.

OpenStudy (anonymous):

Well if you have like terms, you can add the numerators together. You would have 5/3x^2 + 9x/3x^2 = (5+9x)/3x^2

OpenStudy (notamathgenius):

\[\frac{ 5 }{ 3 }x ^{2} + \frac{ 9 }{ 3 }\]

OpenStudy (notamathgenius):

Is that right?

OpenStudy (anonymous):

Don't think so, it looks like a different problem entirely xD

OpenStudy (notamathgenius):

well idk if its 5 over 3 or 5 divided by 3

OpenStudy (anonymous):

Ohhhh, you know what that might have been a good thing to ask first. But I just assumed it was as I said earlier, since one of the answers also agrees with my answer.

OpenStudy (notamathgenius):

@susienicole34 is it divided by or over 3

OpenStudy (anonymous):

over three.

OpenStudy (anonymous):

And I'm not sure what the difference is between dividing by 3 or over 3. It is (5)/(3x^2) + (9)/(3x)

OpenStudy (notamathgenius):

ok so mine was right

OpenStudy (notamathgenius):

But @smokeydabear she was right

OpenStudy (anonymous):

So then what is the answer?